OpenStudy (anonymous):

If A, B and C are the angles of a triangle, prove that sinBcosC+cosBsinC=sinA

OpenStudy (anonymous):

Plug in an angle for each

OpenStudy (anonymous):

A+B+C=180 A+B=180-C sin(A+B)=sin(180-C) sinBcosC+cosBsinC=sinA

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@sauravshakya could you explain the step between sin(A+B)=sin(180-C) sinBcosC+cosBsinC=sinA So sorry and thannks

OpenStudy (anonymous):

Looks like I did a simple mistake, A+B+C=180 B+C=180-A sin(B+C)=sin(180-A) sinBcosC+cosBsinC=sinA

OpenStudy (anonymous):

Thanks alot!

OpenStudy (anonymous):

welcome